OpenStudy (anonymous):

can someone HHHHHHHHHHHHHEEEEEEEEELLPP MATH911!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

tofu on the case!

OpenStudy (saifoo.khan):

Where are u stuck>

OpenStudy (anonymous):

\[6\sqrt{5000}-\sqrt{288}+\sqrt{32}\]

OpenStudy (anonymous):

does anyone have an answer PLEASE

OpenStudy (anonymous):

Are you looking for an explanation or an answer?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

yes i am

OpenStudy (jamesj):

Notice that 5000 = 5 x 1000 = 5 . 10^3 = 5 . 2^3 . 5^3 = 2^3 . 5^4 hence sqrt(5000) = sqrt( 2^3 . 5^4) = sqrt (2^2) . sqrt(5^4) . sqrt(2) = 2 . 5^2 . sqrt(2) = 50 sqrt(2) Now find similar expressions for sqrt(288) and sqrt(32). I.e., write each of them as something times sqrt(2). Then you can just add and subtract these terms. So what are sqrt(288) and sqrt(32) in simplified terms?

OpenStudy (anonymous):

\[5000=2\times 2500\] so \[\sqrt{5000}=\sqrt{2500\times 2}=\sqrt{2500}\times \sqrt{2}=50\sqrt{2}\]

OpenStudy (anonymous):

Looks like JamesJ got it covered..

OpenStudy (anonymous):

thanks anyway

OpenStudy (anonymous):

similarly \[\sqrt{228}=\sqrt{144\times 2}=12\sqrt{2}\] and \[\sqrt{32}=\sqrt{16\times 2}=4\sqrt{2}\]

OpenStudy (jamesj):

@sat73: it's my sense that students at this level don't have a good feel for these higher perfect squares, so I prefer using prime factorizations; but of course what you're writing is correct.

OpenStudy (anonymous):

put it all together and you get \[300\sqrt{2}-12\sqrt{2}+4\sqrt{2}\] etc \[

OpenStudy (anonymous):

i want to ask another question

OpenStudy (anonymous):

pretty sophisticated math for fist grade

OpenStudy (anonymous):

*first

OpenStudy (anonymous):

@jamesj you may be right, but i have noticed that it is usually easier for people to see a perfect square like 144 and 16 than it is to factor a number into the product of primes. at least for smallish numbers

OpenStudy (anonymous):

 Find the area and perimeter of the rectangle. Write each answer in simplest form.